Find the measures of two angles, one positive and one negative, that are coterminal with pi divided by two.

three pi divided by two; negative pi divided by two
pi divided by two + 360°; pi divided by two - 360°
five pi divided by two; negative pi divided by two
five pi divided by two; negative three pi divided by two
1
Expert's answer
2019-11-11T09:10:17-0500

By definition, the angles coterminal with "\alpha" are: "\alpha +360\degree" and "\alpha -360\degree" .(or "\alpha +2\pi" and "\alpha -2\pi)"

  1. "\dfrac{\pi}{2} + 2\pi = \dfrac{5\pi}{2}; \quad \dfrac{\pi}{2} - 2\pi = -\dfrac{3\pi}{2}"
  2. "\dfrac{3\pi}{2} + 2\pi = \dfrac{7\pi}{2}; \quad \dfrac{3\pi}{2} - 2\pi = -\dfrac{\pi}{2}"
  3. "-\dfrac{\pi}{2} + 2\pi = \dfrac{3\pi}{2}; \quad -\dfrac{\pi}{2} - 2\pi = -\dfrac{5\pi}{2}"
  4. "\dfrac{\pi}{2} +2\pi+ 2\pi = \dfrac{9\pi}{2}; \quad \dfrac{\pi}{2} +2\pi - 2\pi = \dfrac{\pi}{2}"
  5. "\dfrac{\pi}{2} -2\pi+ 2\pi = \dfrac{\pi}{2}; \quad \dfrac{\pi}{2} -2\pi - 2\pi = -\dfrac{7\pi}{2}"
  6. "\dfrac{5\pi}{2} + 2\pi = \dfrac{9\pi}{2}; \quad \dfrac{5\pi}{2} - 2\pi = \dfrac{\pi}{2}"
  7. "-\dfrac{\pi}{2} + 2\pi = \dfrac{3\pi}{2}; \quad -\dfrac{\pi}{2} - 2\pi = -\dfrac{5\pi}{2}"
  8. "-\dfrac{3\pi}{2} + 2\pi = \dfrac{\pi}{2}; \quad - \dfrac{3\pi}{2} - 2\pi = -\dfrac{7 \pi}{2}"